if i go 7 miles at 15 miles an hour how long would take
Found 2 solutions by nerdybill, ewatrrr:
Answer by nerdybill(7384)

if i go 7 miles at 15 miles an hour how long would take
applying the distance formula:
d = rt
7 = 15t
7/15 = t
.47 hours = t
or
28 minutes = t

Answer by ewatrrr(24785)

 
Hi,
How long would it take to travel 7mi, traveling at a rate of 15mph
D = r*t or t = D/r
t = 7mi/15mph = .4667 hrs